• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Yi Qiang

    Log entry time 12:01:46 on November 09, 2008

    Entry number 247755

    This entry is a followup to: 247744

    keyword=A huge drop of left VDC effiencicy

    Since run 3108, the lower side of U1 and V1 dropped to about 80% as seen in Chen's halog. However, the decreas of the wire hits are spotted in U1 and U2. Nothing got changed except for stopping/starting a Happex run. Checked High Voltage and Gas flow, all of them seem OK. Beam current, BPM are also fine. The dip of the low efficiency seems moving from low end to high end now, see latest replay from run 3110. Maybe the delay effect of the gas change early this morning?????




    Run# 3107



    Run# 3109



    Run# 3110